OpenStudy (anonymous):

A probability distribution has a mean of 30 and a standard deviation of 2. Use Chebychev's inequality to find a bound on the probability that an outcome of the experiment lies between the following. (Enter your answers to two decimal places.) (a) 25 and 35 (b) 20 and 40

OpenStudy (anonymous):

Chebyshev's inequality says that \[P(|Y-\mu|<k\sigma)\ge1-\frac{1}{k^2}\] where \(\mu\) is the mean of the distribution of \(Y\) and \(\sigma\) is the std. deviation. For part (a), the desired interval is \(25<Y<35\), which can be written as \[-5+30<Y<5+30\\ -5<Y-30<5\\ |Y-\mu|=|Y-30|<5\] Finding \(k\) is easy enough: \[|Y-\mu|<k\sigma~~\Rightarrow~~|Y-30|<2k=5~~\Rightarrow~~k=\frac{5}{2}\] This means that the probability \(Y\) is less than \(2.5\sigma\) from \(\mu\) is \[P\left(|Y-30|<2\cdot\frac{5}{2}\right)\ge\color{red}{1-\frac{1}{(5/2)^2}}=\cdots\] The red part is your desired bound.

OpenStudy (anonymous):

0.84

OpenStudy (anonymous):

Right. Try to set up the second part. It's done very similarly.

OpenStudy (anonymous):

I don't get it though... wouldn't it be the same??

OpenStudy (anonymous):

not the same number of std. devs.

OpenStudy (anonymous):

how many std devs from 20 to 30?

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

no, how much is 1 std dev? think of it as a yard stick.

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

hello sorry I am really confused Im looking in the book again

OpenStudy (anonymous):

how much is 1 std dev? look in the problem statement.

OpenStudy (anonymous):